One sample has n = 4 scores and M = 10. A second sample has n = 6 scores and M = 5. If the two samples are combined, then what is the mean for the combined sample?


1
Expert's answer
2021-11-03T18:08:42-0400

First sample:

"n_1=4, M_1=10, \sum_ix_i=4(10)=40"

Second sample:

"n_2=6, M_2=5, \sum_iy_i=6(5)=30"

Combined sample:


"n_{comb}=n_1+n_2=4+6=10"

"\sum_iu_i=\sum_ix_i+\sum_iy_i=40+30=70"

"M_{comb}=\dfrac{\sum_iu_i}{n_{comb}}=\dfrac{70}{10}=7"

The mean for the combined sample is "M_{comb}=7."